# MixLLM: LLM Quantization with Global Mixed-precision between Output-features and Highly-efficient System Design

### v1

@@ Abstract
Quantization has become one of the most effective methodologies to compress LLMs into smaller size. However, the existing quantization solutions still show limitations of either non-negligible accuracy drop or system inefficiency. In this paper, we make a comprehensive analysis of the general quantization principles on their effect to the triangle of accuracy, memory consumption and system efficiency. We propose MixLLM that explores the new optimization space of mixed-precision quantization between output feature based on the insight that different output feature matter differently in the model. MixLLM identifies the output feature with high salience in the global view rather than within each single layer, effectively assigning the larger bit-width to output features that need it most to achieve good accuracy with low memory consumption. We present the sweet spot of quantization configuration of algorithm-system co-design that leads to high accuracy and system efficiency. To address the system challenge, we design the two-step dequantization to make use of the int8 Tensor Core easily and fast data type conversion to reduce dequantization overhead significantly, and present the software pipeline to overlap the memory access, dequantization and the MatMul to the best. Extensive experiments show that with only 10% more bits, the PPL increasement can be reduced from about 0.5 in SOTA to within 0.2 for Llama 3.1 70B, while on average MMLU-Pro improves by 0.93 over the SOTA of three popular models. In addition to its superior accuracy, MixLLM also achieves state-of-the-art system efficiency.
@@ 1 Introduction
Large language models (LLMs) [7, 32] have shown remarkable performance on various tasks. But their large memory consumption and massive computation cost have become an obstacle for the efficient deployment [43, 44]. Quantization has become one of the most sufficient solution to compress LLMs into smaller size [14, 26, 45, 46], by representing the weight or activation with smaller bit-width. However, the existing quantization solutions still show limitations of either non-negligible accuracy drop or system inefficiency.
There is a triangle of characteristics for efficient LLM quantization: accuracy, memory consumption of parameters, and system efficiency of execution, which we call effectiveness triangle of quantization. The existing quantization solutions have different focus and trade-off in the triangle:
  • The weight-only methodologies target to solve the memory consumption problem, and can speedup the small-batched decoding execution that faces the memory-wall problem [43, 21]. But their accuracy drop of 4-bit quantization can be a challenge for the production workloads sensitive to accuracy, which becomes more serious in the new models with higher information density like Llama 3 [32], as illustrated in recent studies [42, 22]. Besides, the weight-only method can lead to system performance drop for large-batched workloads (e.g., the SOTA W4A16 kernel only achieves 83% performance of its float16 counterpart at batch size 512 with hidden size 4096, shown in Fig.5).
  • The weight-activation quantization represents the activation with low-bit values along with the weights, potentially lead to higher system efficiency. But it can lead to larger accuracy drop than the weight-only method as the activation is usually harder to quantize [50, 3, 27]. Besides, it introduces more dequantization overhead for the activation that can hurt the system efficiency. The transformation optimizations in some works can make the system efficiency even worse.
  • Outlier separation and mixed-precision technologies emerge to improve the accuracy of low-bit quantization by either excluding the unstructured high-salience weights from quantization [11, 21] or assigning larger bit-width for the quantization of structured high-salience weights [50]. The former shows system efficiency problem due to the low efficiency of half precision (i.e., float16/bfloat16) sparse tensor processing. The state-of-the-art mixed-precision solution [50] aims for low-bit quantization but shows non-negligible accuracy drop, even inferior to the 4-bit weight-only quantization.
@@
Contributions. In this paper, we provide an extensive analysis of the general quantization principles. To address the limitations of the previous works and cover the three characteristics in the effectiveness triangle, we propose MixLLM, which makes the following contributions:
@@
Extensive evaluation shows that, with only 10% of 8-bits (i.e., W4.4A8), MixLLM outperforms all the existing 4-bit quantization algorithms while achieving the state-of-the-art system efficiency.
@@ 2 Background, Related Work, and Discussion
2.1 Background of Quantization
The quantizaiton maps the tensor $X$ into the target range with smaller bit-width representation through affine transformation: $X_{q}=clamp(\lfloor\frac{X}{s}\rceil+z,range)$, where $s$ is the scale and $z$ is the zero point. The value can be recovered (i.e., dequantization) through: $X^{\prime}=(X_{q}-z)\times s$. $X^{\prime}$ is pushed to the discrete chunks rather than recovered to the original value, thus has accuracy loss. The bit-width is essential for the accuracy of quantization as it determines the number of chunks for the quantized values ($2^{bit\_width}$). Take an example, enlarging the bit-width from 4 to 5 can double the number of chunks, so that the 5-bit RTN quantization can easily beat the 4-bit quantizations of advanced techniques (Tab.1).
@@ 2.1 Background of Quantization
The scale and zero point can be calculated from the whole channel/token vector or a small group within the channel/token, the former is called per-channel/token quantization and the latter is group-wise quantization. The group-wise scheme results in smaller accuracy loss due to the smaller chunk scale, but requires more complex GPU kernel design. The symmetric quantization uses 0 as the zero point value, which simplifies the computations ($X_{q}=clamp(\lfloor\frac{W}{s}\rceil,range)$, $X^{\prime}=X_{q}\times s$) and enables many works to design the per-channel/per-token quantized kernels by multiplying the scales at the epilogue of the whole MatMul (matrix multiplication) for dequantization [45, 40]. However, it leads to larger loss than the asymmetric one as the data distribution can be usually asymmetric, especially for smaller bit-widths like 4-bit.
2.2 Related Works and Discussion of General Quantization Principles
@@
This paper mainly focuses on post-training quantization (PTQ).
@@
Systems that affect the quantization requirement. The continuous batching technology [47] enables to batch the decoding tasks from different requests together to enlarge the batch dimension of MatMul during LLM inference. The chunked-prefill method [19, 2, 51] advances the continuous batching by merging the prefill and decoding tasks into the same batch, further enlarging the MatMul shapes. These technologies pushes many LLM jobs to become compute-bound and motivate the demand to reduce computation.
@@
Weight-only quantization and its limitation. There emerges a wide range of technologies to improve the accuracy of weight-only quantization. GPTQ [14] advances OBC [13] on OBS-based [18] weight compensation with blocked updating and reordering. AWQ [26] proposes to scale the weight according to the characteristic of activation. OminiQuant [37]) proposes the learnable scaling and weight clipping factors. SpQR [11], SqueezeLLM [21] and OWQ [25] separate the outliers from the quantiation and with half precision. QuiP [8] aims to achieve extreme low-bit quantization with incoherence processing. ZeroQuant(4+2) [42] aims to improve accuracy with medium-sized FP6 quantization.
@@
The weight-only quantization does not reduce the computation but introduces the extra dequantization operations. The low-bit weight will be dequantized to float16 to execute the MatMul in float16 datatype. The current weight-only quantization faces two challenges: 1) From the accuracy aspect, there is still an accuracy gap between the 4-bit quantization and the float16 model, especially for many real business scenarios sensitive to the small accuracy drop, as discussed in the recent works [42, 44]. 2) It can lead to system efficiency drop on busy servers as the recent LLM inference serving systems will usually batch the processing of different requests together on the server and form large MatMuls. The large MatMuls are compute-bound and will suffer from the dequantization overhead [27].
@@
Weight-activation quantization and the challenges. The weight-activation quantization helps to make use of the low-bit computing unit. LLM.int8() [10] observes the activation outlier problem and separates outliers from quantization with half precision. ZeroQuant [46] proposes the per-token activation quantization and group-wise weight quantization. SmoothQuant [45] addresses the activation outlier problem through smoothing, and AffineQuant [29] proposes the general affine transformation for quantization. RPTQ [48] reorders the channels to cluster similar scaled values together. SpinQuant [28] and QuaRot [3] leverages matrix rotation properties to alleviate the outlier phenomenon. Atom [50] uses the mixed-precision between input features to improve accuracy of 4-bit activation quantization. QoQ [27] is a holistic quantization solution with progressive group quantization, attention smoothing, rotation, and channel reordering.
@@
Even though the weight-activation quantization has the advantage of reduced MatMul computation (i.e., MatMul in smaller bit-width to make use of the smaller bit-width computing unit with higher computation throughput [^2] The extra dynamic activation quantization kernel can be fused into other operators with very little system cost [50], thus we only discuss the MatMul itself.), it faces the challenge of accuracy drop caused by the activation quantization, especially that the activation is usually harder to quantize than the weight. The SOTA low-bit weight-activation solutions [3, 28, 27] still have a gap to the 4-bit weight only quantization.
@@
Beside the accuracy drop, the activation quantization will introduce more dequantization overhead than the weight-only one, which makes it challenging to design efficient GPU kernels. When enabling the asymmetric quantization, the result of $(X_{q}-z)$ may exceed the range of the bit-width of $X_{q}$, making it hard to use the corresponding Tensor Core computation. Systems like Atom [50] thus avoid using the asymmetric quantization, with the cost of larger accuracy drop. The group-wise quantization requires fine-grained integer-to-float (I2F) conversion to apply per-group scales. However, the I2F instruction is more expensive than the common computation instructions on the GPU [1] and can lead to large system performance drop ($>10\%$ drop in our practice). Besides, the throughput of Tensor Core is much higher than that of SIMT Cores, 624 TOPS of int8 Tensor Core vs. 19.5 TFLOPS/TOPS of FP32/INT32 SIMT Cores. There still lacks a well designed software pipeline to overlap the Tensor Core computation and SIMT Core based dequantization in the existing works while achieving a high accuracy.
In general, the existing solutions focus on partial of the effectiveness triangle, but cannot cover all of them well. MixLLM is orthogonal to the above works by exploring the mixed-precision between output features with global salience identification, and the co-designed quantization decision and GPU kernels.
@@ 3.1 Quantization Design and Decision in MixLLM
To cover the three aspects of the effectiveness triangle simultaneously, we make the following design and decision of weight and activation quantization according to the analysis in Sec.2.2.
3.1.1 Mixed-precision between output features of weight, with global salience identification.
It is known that different elements of the weight show different salience to the network’s loss when being quantized [21, 11]. The outlier separation method can improve the accuracy by using float16 to store the high-salience elements, but can suffer from the inefficient sparse MatMul. We observe that the elements with high salience tend to show distribution along the output channels for most of the linear layers in many LLMs. Based on this observation, we can assign larger bit-width to the output channels of high salience, and smaller bit-with to the others, forming structured mixed-precision quantization. Through the experiments, we get the same conclusion with the existing works [21, 11] that there is only a small set of elements with high salience contributing significantly to the model’s accuracy drop. Thus we only need to assign the large bit-width to a small portion of the output channels to achieve good accuracy and retain a small memory consumption at the same time.
The structured mixed-precision between different output channels can be friendly to the system efficiency and kernel development, due to the nature that different output features are disjoint in the MatMul and the computation of them are different sub-problems. Fig.1 shows how the linear layer computes with the mixed-precision between output features. It divides the linear into independent sub-problems, and finally gathers the output of the sub-problems together to form the final result. This optimization space is orthogonal to the existing quantization optimizations, e.g., GPTQ [14], and can be applied together with them.
One critical problem is how to identify the high-salience output channels in the model. The fixed threshold [11] or the fixed number/ratio [50, 25] of high salience elements computed by the local loss of layers can be sub-optimal to the end-to-end model, as different layers can show different importance to the model’s final output [16, 30, 12]. A high salience channel w.r.t. a layer may not be a high salience channel of the end-to-end model. In MixLLM, we compute the high salience channels globally according to their impact to the model’s final loss (Sec.3.2). As a result, different layers will have different number of high salience channels. Fig.2 shows the distribution of the top 10% high-salient out features in Llama 3.1 8B, showing huge difference in different linear layers.
@@
Note that this design is different from the mixed-precision in Atom [50] from two aspects. 1) MixLLM first addresses the problem of identifying the high-salience channels globally rather than locally. 2) MixLLM applies the mixed-precision between output features rather than input features, which is more system performant and algorithm flexible [^3] One example is that, the outlier number in each MatMul should be a multiplier of the corresponding tiling size of kernel design to achieve a good system efficiency in Atom, which limits the flexibility of algorithm. as the output features are disjoint naturally.
3.1.2 Quantization decision with algorithm-system consideration: 8-bit symmetric activation and 4-bit asymmetric weight quantization in group-wise manner.
MixLLM makes the same decision with QoQ [27] on activation quantization to use 8-bit, as the 4-bit activation can lead to a large accuracy drop but does not lead to significant system efficiency improvement as MatMul execution tends to be bound more on the larger weight tensor rather than the smaller activation tensor. It can be partially indicated from the compute intensity of the linear layer. Given token number $M$ and in/out features $K$/$N$, the compute intensity $I=\frac{2MNK}{MKB_{act}+KNB{weight}}$. $B_{act}$ and $B_{weight}$ are the bytes per element of activation and weight. Given $M=512$ and $N=K=4096$, reducing $B_{weight}$ from 8 to 4 will results in an 80% increasement of $I$, while reducing $B_{act}$ from 8 to 4 will only achieve 5.88% increasement.
Instead of using per-token activation quantization, MixLLM uses group-wise RTN method. On the one hand, Tab.1 shows that simple group-wise RTN quantization performs better than token-wise smoothing method. On the other hand, the weight is already group-wise in MixLLM, and the group-wise activation does not lead to significant more dequantization overhead in the system. We observe symmetric quantization is enough for the 8-bit activation (refer to MixLLM W8A8 in Tab.1), while asymmetric can be essential for the 4-bit weight. The group-wise method with asymmetric can lead to a difficulty for the kernel to make use int8 Tensor Core, for which QoQ [27] introduces the two-step quantization method. Instead, we design a two-step dequantization method with the property of the mix of symmetric and asymmetric (Sec.3.3).

### v2

@@ Abstract
Quantization has become one of the most effective methodologies to compress LLMs into smaller size. However, the existing quantization solutions still show limitations of either non-negligible accuracy drop or low system efficiency. In this paper, we propose MixLLM that explores the optimization space of mixed-precision quantization between output features, based on the insight that different features matter differently in the model. MixLLM identifies the important output features in the global view rather than within each single layer, effectively assigning larger bit-width to output features that need it the most to achieve high accuracy and low memory usage. We present the sweet spot of quantization configuration of algorithm-system co-design with high accuracy and system efficiency. To address the system challenge, we design the two-step dequantization to make use of the Tensor Core easily and fast data type conversion to reduce dequantization overhead, and present the software pipeline to overlap the memory access, dequantization and the MatMul to the best. Extensive experiments show that with only 10% more bits, the perplexity increase can be reduced from about 0.5 in SOTA to within 0.2 for Llama 3.1 70B, while MMLU-Pro loss can be reduced from 1.92 to 0.99 over the SOTA of three popular models. Besides its superior accuracy, MixLLM also achieves state-of-the-art system efficiency. Code is released at https://github.com/microsoft/MixLLM.
footnotetext: 1Anonymous Institution, Anonymous City, Anonymous Region, Anonymous Country. Correspondence to: Anonymous Author <anon.email@domain.com>.
Preliminary work. Under review by the Machine Learning and Systems (MLSys) Conference. Do not distribute.
@@ 1 Introduction
Large language models (LLMs) Bubeck et al. (2023); Meta (Cited 2024) have shown remarkable performance on various tasks. But their large memory consumption and massive computation cost have become an obstacle for the efficient deployment Xia et al. (2023; 2024). Quantization has become one of the most effective solution to compress LLMs into smaller size Frantar et al. (2022); Lin et al. (2024); Xiao et al. (2023); Yao et al. (2022), by representing the weight or activation with smaller bit-width. However, the existing quantization solutions still show limitations of either non-negligible accuracy drop or system inefficiency.
There is a triangle of characteristics for LLM quantization: accuracy, memory consumption of parameters, and system efficiency of execution, The existing quantization solutions have different focus and trade-off in the triangle: 1) The weight-only method targets to solve the memory consumption problem, and can speedup the small-batched decoding execution that faces the memory-wall problem Xia et al. (2023); Kim et al. (2024). But their accuracy drop of 4-bit quantization can be a challenge for the production workloads sensitive to accuracy, as illustrated in recent studies Wu et al. (2023); Kumar et al. (2024). Besides, the weight-only method can lead to system performance drop for large-batched workloads due to the dequantization overhead. 2) The weight-activation quantization represents the activation with low-bit values along with the weights, potentially lead to higher system efficiency. But it can lead to larger accuracy drop than the weight-only method as the activation is usually harder to quantize Zhao et al. (2024); Ashkboos et al. (2024); Lin et al. (2025). Besides, it introduces more dequantization overhead for the activation that can hurt the system efficiency.
@@
Contributions. In this paper, we provide an extensive analysis of the general quantization principles, and propose MixLLM with the following contributions:
@@
Extensive evaluation shows that, with 10% 8-bit and 90% 4-bit (i.e., W4.4A8), MixLLM significantly outperforms SOTA quantization algorithms while achieving the state-of-the-art system efficiency on both A100 and H100 GPUs.
@@ 2 Background, Related Work, and Discussion
Quantization maps the tensor $X$ into the target range with smaller bit-width through $X_{q}=clamp(\lfloor\frac{X}{s}\rceil+z,range)$, where $s$ is the scale and $z$ is the zero point. The scale and zero point can be calculated from the whole channel/token vector or a smaller group. The group-wise scheme results in higher accuracy but requires more complex GPU kernel design. The symmetric quantization uses 0 as the zero point, which simplifies the computations and enables many works to design the per-channel/per-token quantized kernels by multiplying the scales at the epilogue of the whole MatMul (matrix multiplication) for dequantization Xiao et al. (2023); TensorRT-LLM (Cited 2024). However, it usually leads to larger loss than the asymmetric one, especially for smaller bit-widths like 4-bit.
2.1 Related Work and Discussion
@@
This paper focuses on pure post-training quantization (PTQ), without the training on the weight (i.e., QAT) or the quantization parameters (e.g., training the rotation matrix like in SpinQuant Liu et al. (2024)).
@@
Systems that affect the quantization requirement. The continuous batching technology Yu et al. (2022) enables to batch the decoding tasks from different requests together to enlarge the batch dimension of MatMul during LLM inference. The chunked-prefill method Holmes et al. (2024); Agrawal et al. (2024); Zheng et al. (2024) advances the continuous batching by merging the prefill and decoding tasks into the same batch, further enlarging the MatMul shapes. These technologies pushes many LLM jobs to become compute-bound and motivate the demand to reduce computation.
@@
Weight-only quantization and its limitation. There emerges a wide range of technologies to improve the accuracy of weight-only quantization. GPTQ Frantar et al. (2022) advances OBC Frantar and Alistarh (2022) on OBS-based Hassibi et al. (1993) weight compensation with blocked updating and reordering. AWQ Lin et al. (2024) proposes to scale the weight according to the characteristic of activation. OmniQuant Shao et al. (2024) proposes the learnable scaling and weight clipping factors. SpQR Dettmers et al. (2024), SqueezeLLM Kim et al. (2024) and OWQ Lee et al. (2024) separate the outliers from the quantiation and with half precision. QuIP Chee et al. (2023) and QuIP# Tseng et al. (2024a) aim to achieve extremely low-bit quantization with incoherence processing, QTIP Tseng et al. (2024b) also leverages incoherence processing. AQLM Egiazarian et al. (2024) and PV-Tuning Malinovskii et al. (2024) target extremely low-bit quantization. SliM-LLM Huang et al. (2024) focuses on extremely low-bit quantization using intra-layer mixed-precision. ZeroQuant(4+2) Wu et al. (2023) and Quant-LLM Xia et al. (2024) aim to improve accuracy with FP6 quantization.
@@
The weight-only quantization does not reduce the computation but introduces the extra dequantization operations, as it requires to dequantize the weight into float16 for execution. The current weight-only quantization faces two challenges: 1) From the accuracy aspect, there is still an accuracy gap between the 4-bit quantization and the float16 model, especially for many real business scenarios sensitive to the small accuracy drop, as discussed in the recent works Wu et al. (2023); Xia et al. (2024). 2) It can lead to system efficiency drop on busy servers as the recent LLM inference serving systems will usually batch the processing of different requests together on the server and form large MatMuls. The large MatMuls are compute-bound and will suffer from the dequantization overhead Lin et al. (2025).
@@
Weight-activation quantization and the challenges. The weight-activation quantization helps to make use of the low-bit computing unit. LLM.int8() Dettmers et al. (2022) observes the activation outlier problem and separates outliers from quantization with half precision. ZeroQuant Yao et al. (2022) proposes the per-token activation quantization and group-wise weight quantization. SmoothQuant Xiao et al. (2023) addresses the activation outlier problem through smoothing, and AffineQuant Ma et al. (2024) proposes the general affine transformation for quantization. RPTQ Yuan et al. (2023) reorders the channels to cluster similar scaled values together. SpinQuant Liu et al. (2024) and QuaRot Ashkboos et al. (2024) leverages matrix rotation properties to alleviate the outlier phenomenon. Atom Zhao et al. (2024) uses the mixed-precision between input features to improve accuracy of 4-bit activation quantization. QServe Lin et al. (2025) is a holistic solution of W4A8 quantization.
@@
Even though the weight-activation quantization has the advantage of reduced MatMul computation (i.e., smaller bit-width computation with higher throughput), it faces the challenge of accuracy drop caused by the activation quantization. The SOTA low-bit weight-activation solutions Ashkboos et al. (2024); Liu et al. (2024); Lin et al. (2025) still have a gap to the 4-bit weight only quantization. Besides the accuracy drop, the activation quantization will introduce more dequantization overhead than the weight-only one, which makes it challenging to design efficient GPU kernels.
@@
When enabling the asymmetric quantization, the result of $(X_{q}-z)$ may exceed the range of the bit-width of $X_{q}$, making it hard to use the corresponding Tensor Core computation. Systems like Atom Zhao et al. (2024) thus avoid using the asymmetric quantization, with the cost of larger accuracy drop. The integer quantization requires integer-to-float (I2F) conversion to apply scales. However, the I2F instruction is more expensive than the common computations on modern GPUs Abdelkhalik et al. (2022) and can lead to large system performance drop for group-wise quantization ($10\%$ drop in our practice). Besides, the throughput of Tensor Core is much higher than SIMT Cores, 624 TOPS of int8 Tensor Core vs. 19.5 TFLOPS/TOPS of FP32/INT32 SIMT Cores on A100 GPU. There lacks a well designed software pipeline to overlap the Tensor Core computation and SIMT Core based dequantization for the group-wise and asymmetric weight-activation quantization.
Performance challenge of the float16 outlier separation Outlier separation with half precision works to improve the accuracy while using small bit-width for the non-sensitive weights Kim et al. (2024); Dettmers et al. (2024), by separating the outliers into an extra sparse tensor in float16. However, it is hard to achieve the peak performance due to the inefficiency of the sparse computation on the GPU, especially when the batch size is large and the linear layer becomes compute-bounded. (As discussed in Flash-LLM Xia et al. (2023), the hardware utilization can be lower than 10% for the sparse MatMul, while its dense counterpart can usually achieve more than 60%.) This is because the unstructured tensor computation cannot make use the fast Tensor Core easily, but has to use the SIMT Core in float16 for computation and float32 for accumulation [^1] Flash-LLM Xia et al. (2023) optimizes the unstructured sparse MatMul, but can only speedup the small-batched scenarios.. Moreover, sparse computing makes it more difficult to fully utilize the hardware due to the non-continuous memory pattern and the extra index computation.
@@ 3.1 Quantization Design and Decision in MixLLM
We make the following design and decision to optimize the quantization algorithm and system.
3.1.1 Global Mixed-precision
Different elements of the weight show different salience to the network’s loss when being quantized Kim et al. (2024); Dettmers et al. (2024). The outlier separation method can improve the accuracy by using float16 for high-salience elements, but can suffer from the inefficient sparse MatMul. We observe that the elements with high salience tend to show distribution along the output channels for most of the linear layers in many LLMs. Based on this observation, we can assign larger bit-width to the output channels of high salience, and smaller bit-width to the others. Through the experiments, we get the same conclusion with the existing works Kim et al. (2024); Dettmers et al. (2024) that there is only a small set of elements with high salience to quantization. Thus we only need to assign the large bit-width to a small portion of the output channels to achieve good accuracy while retaining a small memory consumption.
The structured mixed-precision between output channels can be friendly to the system efficiency and kernel development, due to the nature that different output features are disjoint in the MatMul and the computation of them are different sub-problems. Fig.1 illustrates the mixed-precision quantization and execution in MixLLM. It divides the linear into independent sub-problems, and finally gathers the output of the sub-problems together to form the result. This optimization space is orthogonal to the existing quantization methodologies and can be applied together with them.
@@ 3.1.1 Global Mixed-precision
One critical problem is how to identify the high-salience output channels in the model. The fixed threshold Dettmers et al. (2024) or the fixed number/ratio Zhao et al. (2024); Lee et al. (2024) of high salience elements computed by the local loss of layers can be sub-optimal to the end-to-end model, as different layers show different importance to the model’s final output Gromov et al. (2024); Men et al. (2024); Dong et al. (2019). A high salience channel w.r.t. a layer may not be a high salience channel of the end-to-end model. In MixLLM, we compute the high salience channels globally according to their impact to the model’s final loss (Sec.3.2). As a result, different layers will have different number of high salience channels. Fig.2 shows the distribution of the top 10% high-salient out features in Llama 3.1 8B, showing huge difference between different layers.
Note that this design is different from the mixed-precision in Atom Zhao et al. (2024) from two aspects. 1) MixLLM first addresses the problem of identifying the high-salience channels globally rather than locally. 2) MixLLM applies the mixed-precision between output features rather than input features, which is more system/algorithm flexible as the output features are disjoint naturally. It also differs from SliM-LLM Huang et al. (2024) which also only considers the local loss to determine the mixed precision and does not focus on the system performance problem.
3.1.2 Quantization Decision
MixLLM makes the same decision with QServe Lin et al. (2025) on activation quantization to use 8-bit, as the 4-bit activation can lead to a large accuracy drop but does not lead to significant system efficiency improvement as MatMul execution tends to be bound more on the larger weight tensor rather than the smaller activation tensor according to the compute intensity $I=\frac{2MNK}{MKB_{act}+KNB{weight}}$ ($M$ is the token number, KK/NN are the in/out features, and $B_{act}$ and $B_{weight}$ are the bytes per element of activation and weight).
Instead of using per-token activation quantization, MixLLM uses group-wise RTN method. On the one hand, Tab.1 shows that simple group-wise RTN quantization outperforms token-wise smoothing method. On the other hand, the weight is already group-wise in MixLLM, and the group-wise activation does not lead to significantly more dequantization overhead in the system. We observe symmetric quantization is enough for the 8-bit activation (refer to MixLLM W8A8 in Tab.1), while asymmetric is essential for the 4-bit weight. The group-wise method with asymmetric can lead to difficulty for the kernel to make use int8 Tensor Core, for which we design a two-step dequantization method with the property of the mix of symmetric and asymmetric (Sec.3.3).